# Building Enterprise Realtime Voice Agents from Scratch: A Technical Tutorial

### 05 Mar 2026 v1

@@ Abstract
We present a technical tutorial for building enterprise-grade realtime voice agents from first principles. While over 25 open-source speech-to-speech models and numerous voice agent frameworks exist, no single resource explains the complete pipeline from individual components to a working streaming voice agent with function calling capabilities. Through systematic investigation, we find that (1) native speech-to-speech models like Qwen2.5-Omni, while capable of high-quality audio generation, are too slow for realtime interaction ($\sim$13s time-to-first-audio); (2) the industry-standard approach uses a cascaded streaming pipeline: STT $\rightarrow$ LLM $\rightarrow$ TTS, where each component streams its output to the next; and (3) the key to “realtime” is not any single fast model but rather streaming and pipelining across components. We build a complete voice agent using Deepgram (streaming STT), vLLM-served LLMs with function calling (streaming text generation), and ElevenLabs (streaming TTS), achieving a measured P50 time-to-first-audio of 947ms (best case 729ms) with cloud LLM APIs, and comparable latency with self-hosted vLLM on NVIDIA A10G GPU. We release the full codebase as a tutorial with working, tested code for every component.
@@ 1 Introduction
  • 25+ open-source speech-to-speech models exist (Moshi [moshi2024], Qwen2.5-Omni [qwen25omni2025], GLM-4-Voice [glm4voice2024], Kimi-Audio [kimiaudio2025], Step-Audio [stepaudio2025], LLaMA-Omni [llamaomni2024], Mini-Omni [miniomni2024], Freeze-Omni [freezeomni2024], among others [wavchat2024, spiritlm2024]), but none ships with a tutorial on building a realtime voice agent.
@@
  • No single codebase builds a complete streaming enterprise voice agent from scratch, explaining every component.
@@
    An empirical evaluation of native speech-to-speech (Qwen2.5-Omni) vs. cascaded streaming pipelines, demonstrating why the latter is necessary for realtime (Section˜3).
@@ 2.1 Speech-to-Speech Models
A text LLM is augmented with a speech encoder and speech decoder, jointly trained. Speech goes in and out, but “thinking” happens in text space. This is the largest category, including Qwen2.5-Omni [qwen25omni2025] (Thinker-Talker architecture), GLM-4-Voice [glm4voice2024], Kimi-Audio [kimiaudio2025], Step-Audio [stepaudio2025], LLaMA-Omni [llamaomni2024], Mini-Omni [miniomni2024], and Freeze-Omni [freezeomni2024].
@@
Separate models handle speech recognition, language understanding, and speech synthesis. This is the architecture used by all production voice agents (Vapi, Retell, Bland.ai, and systems built with Pipecat/LiveKit). While less elegant, it offers maximum flexibility and, crucially, supports function calling.
Table˜1 summarizes the key models. A critical observation: none of the Level 1 or Level 2 models support function calling, which is essential for enterprise voice agents. This fundamentally constrains the architecture choice.
Table 1: Representative speech-to-speech models with open weights. “FC” = function calling support. “FD” = full-duplex. “Streaming” = supports incremental audio output.
@@ 2.2 Voice Agent Frameworks
3 Why Native Speech-to-Speech Is Not (Yet) Realtime
We conducted an empirical evaluation of Qwen2.5-Omni-7B [qwen25omni2025] to determine whether native speech-to-speech models can serve as the foundation for realtime voice agents.
3.1 Experimental Setup
Qwen2.5-Omni uses a “Thinker-Talker” architecture: the Thinker (a text LLM) generates text tokens, and the Talker (a DiT-based audio decoder) generates speech conditioned on the Thinker’s hidden states. We loaded the 7B model in BF16 on a single NVIDIA A10G GPU with SDPA attention.
Note: transformers version matters.
We discovered that transformers>=5.0 produces low-quality, noisy audio with Qwen2.5-Omni. Only version 4.52.3 (the version recommended by the model card) produces high-quality speech.
3.2 Results
Table 2: Qwen2.5-Omni-7B inference latency on a single NVIDIA A10G GPU. “Text only” disables the Talker; “Text + Audio” runs the full pipeline.
@@ 3.2 Results
As shown in Table˜2, the Talker operates at approximately $0.5\times$ realtime, it takes roughly 2 seconds to generate 1 second of audio. This makes the time-to-first-audio (TTFA) fundamentally too high for realtime interaction:
  • Batch mode: 26.5s TTFA for a 13.8s response.
  • Sentence-level streaming: 13.2s TTFA (generating audio per sentence as text streams).
  • Text-only mode: 0.71s, fast enough, but no audio output.
The Talker is the bottleneck.
The DiT-based vocoder in Qwen2.5-Omni cannot produce partial audio incrementally. Each model.generate() call with return_audio=True blocks until all audio is synthesized.
vLLM cannot help.
vLLM’s Qwen2.5-Omni support is “Thinker only”, where it can accelerate text generation but cannot run the Talker. The vLLM example is literally named only_thinker.py.
Conclusion.
@@ 4.3 Component 2: Streaming LLM (vLLM)
We serve the LLM using vLLM [vllm2023] on an NVIDIA A10G GPU, exposing an OpenAI-compatible API at /v1/chat/completions. Because vLLM implements the OpenAI API specification, the same client code works interchangeably with self-hosted vLLM, the OpenAI API, Azure OpenAI, or any OpenAI-compatible endpoint, simply by changing the base_url and api_key environment variables.
@@ 8 Experimental Results
With self-hosted vLLM (Qwen2.5-7B-Instruct on A10G)
@@
Key findings from Table˜3:
@@
Table 4: Time-to-first-audio comparison.
Approach TTFA Function Calling
Qwen2.5-Omni (batch) $\sim$26,500ms
@@
[…]
9 Practical Notes
  1. 1.
    Transformers version: transformers>=5.0 produces noisy audio with Qwen2.5-Omni. Only ==4.52.3 works correctly.
  2. 2.
    Single GPU loading: device_map=‘‘auto’’ across multiple GPUs degrades audio quality. Use device_map=‘‘cuda:0’’.
  3. 3.
    System prompt sensitivity: Qwen2.5-Omni’s audio generation requires the exact training system prompt.
  4. 4.
    Speech speed control: Prepending “Please speak quickly.” naturally increases speech speed without quality loss.
  5. 5.
    vLLM version: v0.16.0 has a SageMaker dependency that fails outside AWS. v0.8.5 works reliably.
  6. 6.
  7. 7.
  8. 8.
    Empty SSE chunks: Always guard with if not chunk.choices: continue.
  9. 9.
@@ 10 Related Work
The WavChat survey [wavchat2024] provides a comprehensive overview of spoken dialogue models. The field has seen rapid growth since GPT-4o’s voice mode demonstration in May 2024 [gpt4o2024]. Level 1 (truly native) models include Moshi [moshi2024], which achieves $\sim$200ms latency with a 7B dual-stream architecture using the Mimi codec, and is the only model shipping a complete serving stack (Rust+CUDA server, WebSocket protocol, web client). Spirit LM [spiritlm2024] from Meta interleaves speech and text tokens but is research-only. Level 2 (speech-augmented LLMs) models are the largest category: Qwen2.5-Omni [qwen25omni2025] uses a Thinker-Talker architecture with TMRoPE for multimodal alignment; GLM-4-Voice [glm4voice2024] employs a flow-matching decoder for streaming synthesis; Kimi-Audio [kimiaudio2025] uses a 7B model trained on 13M+ hours; Step-Audio [stepaudio2025] scales to 130B parameters; LLaMA-Omni [llamaomni2024] achieves 226ms latency with Llama-3.1-8B; Mini-Omni [miniomni2024] demonstrates “talking while thinking” with simultaneous text and audio generation; and Freeze-Omni [freezeomni2024] keeps the text LLM frozen to prevent catastrophic forgetting. Common across all Level 2 models is the use of a text LLM backbone (predominantly Qwen2.5 [qwen25instruct2024]) with speech encoder/decoder modules. Critically, none of these models support function calling, which limits their applicability to enterprise voice agents.
@@
Unlike the frameworks above which provide ready-made components, our contribution is educational: a from-scratch implementation that explains how each piece works internally. Unlike native S2S models which prioritize end-to-end elegance, we prioritize practical latency and function calling support through the cascaded approach.
@@ 11 Conclusion
    Native S2S models are not yet ready for realtime: Qwen2.5-Omni achieves $\sim$13s TTFA, lacks function calling, and has no incremental audio output. The cascaded approach achieves $\sim$755ms TTFA with full function calling.
@@
    Realtime = streaming + pipelining: The perception of instant response comes from overlapping execution of STT, LLM, and TTS. The sentence buffer is the critical orchestration primitive.
@@
    Sub-1-second TTFA is achievable: With Deepgram ($\sim$337ms), vLLM ($\sim$337ms TTFT), and ElevenLabs ($\sim$220ms TTFB), the measured best-case TTFA is 729ms.

### 17 Mar 2026 v2

@@ Abstract
We present a technical tutorial for building enterprise-grade realtime voice agents from first principles. While end-to-end speech-to-speech models may ultimately provide the best latency for voice agents, fully self-hosted end-to-end solutions are not yet available. We evaluate the closest candidate, Qwen3-Omni, across three configurations: its cloud-only DashScope Realtime API achieves $\sim$702ms audio-to-audio latency with streaming, but is not self-hostable; its local vLLM deployment supports only the Thinker (text generation from audio, 516ms), not the Talker (audio synthesis); and its local Transformers deployment runs the full pipeline but at $\sim$146s—far too slow for realtime. The cascaded streaming pipeline (STT $\rightarrow$ LLM $\rightarrow$ TTS) therefore remains the practical architecture for self-hosted realtime voice agents, and the focus of this tutorial. We build a complete voice agent using Deepgram (streaming STT), vLLM-served LLMs with function calling (streaming text generation), and ElevenLabs (streaming TTS), achieving a measured time-to-first-audio of 755ms (best case 729ms) with full function calling support. We release the full codebase as a 9-chapter progressive tutorial with working, tested code for every component.
@@ 1 Introduction
  • 25+ open-source speech-to-speech models exist (Moshi [moshi2024], Qwen3-Omni [qwen3omni2025], Qwen2.5-Omni [qwen25omni2025], GLM-4-Voice [glm4voice2024], Kimi-Audio [kimiaudio2025], Step-Audio [stepaudio2025], LLaMA-Omni [llamaomni2024], Mini-Omni [miniomni2024], Freeze-Omni [freezeomni2024], among others [wavchat2024, spiritlm2024]), but none ships with a tutorial on building a realtime voice agent.
@@
  • No codebase builds a complete streaming enterprise voice agent from scratch, explaining every component.
@@
    An empirical evaluation of native speech-to-speech models (Qwen2.5-Omni and Qwen3-Omni) vs. cascaded streaming pipelines, demonstrating that while end-to-end models are approaching realtime, a fully self-hosted solution is not yet available (Section˜3).
@@ 2.1 Speech-to-Speech Models
A text LLM is augmented with a speech encoder and speech decoder, jointly trained. Speech goes in and out, but “thinking” happens in text space. This is the largest category, including Qwen3-Omni [qwen3omni2025] (MoE-based Thinker-Talker with multi-codebook audio decoder), Qwen2.5-Omni [qwen25omni2025] (Thinker-Talker architecture), GLM-4-Voice [glm4voice2024], Kimi-Audio [kimiaudio2025], Step-Audio [stepaudio2025], LLaMA-Omni [llamaomni2024], Mini-Omni [miniomni2024], and Freeze-Omni [freezeomni2024]. Notably, Qwen3-Omni is the first Level 2 model to support function calling.
@@
Separate models handle speech recognition, language understanding, and speech synthesis. This is the architecture used by all production voice agents (Vapi, Retell, Bland.ai, and systems built with Pipecat/LiveKit). It offers maximum flexibility and full self-hosted control.
Table˜1 summarizes the key models. Until recently, no Level 1 or Level 2 model supported function calling, which is essential for enterprise voice agents. Qwen3-Omni [qwen3omni2025] changes this: it supports function calling via <tool_call> XML tags, making it the first native S2S model viable for enterprise use.
Table 1: Representative speech-to-speech models with open weights. “FC” = function calling support. “FD” = full-duplex. “Streaming” = supports incremental audio output. Streaming audio via DashScope Realtime API; local vLLM supports text streaming only (Thinker).
@@ 2.2 Voice Agent Frameworks
3 Native Speech-to-Speech: From Qwen2.5-Omni to Qwen3-Omni
We empirically evaluate two generations of native speech-to-speech models to assess whether they can serve as the foundation for realtime voice agents.
3.1 Qwen2.5-Omni-7B (Baseline)
Qwen2.5-Omni [qwen25omni2025] uses a “Thinker-Talker” architecture: the Thinker (a 7B dense text LLM) generates text tokens, and the Talker (a DiT-based audio decoder) generates speech conditioned on the Thinker’s hidden states. We loaded the model in BF16 on a single NVIDIA H200 GPU.
Table 2: Qwen2.5-Omni-7B inference latency on a single NVIDIA H200 GPU.
@@
3.2 Qwen3-Omni-30B-A3B (New)
Qwen3-Omni [qwen3omni2025] introduces several architectural upgrades: (1) a Mixture-of-Experts (MoE) backbone with 30B total parameters but only 3B active per token, (2) a multi-codebook Talker designed for low-latency streaming audio synthesis, and (3) native function calling via <tool_call> XML tags. We evaluate it across three deployment configurations.
Configuration 1: vLLM (Thinker only, self-hosted).
Using vLLM 0.13.0 on a single NVIDIA H200 GPU, we run the Thinker for text generation from audio input. This configuration eliminates the need for a separate STT component.
Configuration 2: Transformers (Thinker + Talker, self-hosted).
Using HuggingFace Transformers with return_audio=True on a single H200. This runs the full end-to-end pipeline but uses the unoptimized transformers backend.
Configuration 3: DashScope Realtime API (cloud).
The DashScope API (qwen3-omni-flash-realtime) provides optimized end-to-end audio-to-audio streaming via WebSocket, with server-side VAD and streaming audio output.
Table 3: Qwen3-Omni inference latency across deployment configurations. TTFT = time-to-first-text, TTFA = time-to-first-audio, FC = function calling.
Configuration TTFT TTFA tok/s FC Audio Out
[…]
DashScope Realtime API 449ms 702ms Partial ✓ (streaming)
Audio-to-text latency (no audio output); still requires external TTS for speech.
3.3 Analysis
vLLM Thinker is production-viable.
At 516ms audio-to-text with 168 tok/s throughput, the Qwen3-Omni Thinker served via vLLM can replace both the STT and LLM stages of a cascaded pipeline, reducing a 3-step pipeline (STT $\rightarrow$ LLM $\rightarrow$ TTS) to a 2-step pipeline (Qwen3-Omni $\rightarrow$ TTS).
Transformers backend is too slow for realtime.
[…]
DashScope API achieves realtime audio-to-audio.
At 702ms average TTFA (685ms best), the DashScope Realtime API demonstrates that Qwen3-Omni’s multi-codebook Talker can achieve realtime streaming audio generation when properly optimized. This is comparable to our cascaded pipeline’s 755ms TTFA. However, this requires the cloud API—no self-hosted solution for the optimized Talker exists yet.
Key takeaway.
Qwen3-Omni represents significant progress over Qwen2.5-Omni: it supports function calling and its multi-codebook Talker is architecturally designed for realtime streaming. However, the optimized Talker is only available through the cloud-hosted DashScope API—there is no self-hosted solution for realtime audio generation. Locally, vLLM can only run the Thinker (text output), and the Transformers backend is far too slow for the full pipeline. Until an optimized, self-hostable Talker serving solution exists (e.g., vLLM-Omni adding Talker support), the cascaded pipeline remains the only viable architecture for fully self-hosted realtime voice agents. This is the motivation for our tutorial.
@@ 4.3 Component 2: Streaming LLM (vLLM)
We serve the LLM using vLLM [vllm2023] on an NVIDIA H200 GPU, exposing an OpenAI-compatible API at /v1/chat/completions. Because vLLM implements the OpenAI API specification, the same client code works interchangeably with self-hosted vLLM, the OpenAI API, Azure OpenAI, or any OpenAI-compatible endpoint, simply by changing the base_url and api_key environment variables.
@@ 8 Experimental Results
With self-hosted vLLM (Qwen2.5-7B-Instruct on H200)
@@
Findings from Table˜4:
@@
Table 5: Time-to-first-audio comparison across all approaches. “Self-hosted” = runs entirely on local GPUs without external API dependencies.
Approach TTFA FC Audio Out Self-hosted
@@
Qwen3-Omni (transformers, full) $\sim$145,694ms
Qwen3-Omni (vLLM Thinker + TTS) $\sim$736ms
Qwen3-Omni (DashScope Realtime) $\sim$702ms Partial
@@
Estimated as vLLM audio-to-text (516ms) + ElevenLabs TTS TTFB (220ms). This 2-step pipeline eliminates the STT stage.
Table˜5 highlights the current state of affairs. While Qwen3-Omni’s DashScope API demonstrates that end-to-end audio-to-audio at $\sim$702ms is architecturally feasible, this requires a cloud API and is not self-hostable. The only fully self-hosted end-to-end option (Transformers) is impractical at $\sim$146s. A hybrid 2-step approach (Qwen3-Omni vLLM Thinker + external TTS) is promising at $\sim$736ms but still depends on external TTS. Our cascaded pipeline achieves $\sim$755ms with full self-hosted control, robust function calling, and component-level flexibility—making it the most practical fully self-hosted architecture today.
@@ 9 Related Work
The WavChat survey [wavchat2024] provides a comprehensive overview of spoken dialogue models. The field has seen rapid growth since GPT-4o’s voice mode demonstration in May 2024 [gpt4o2024]. Level 1 (truly native) models include Moshi [moshi2024], which achieves $\sim$200ms latency with a 7B dual-stream architecture using the Mimi codec, and is the only model shipping a complete serving stack (Rust+CUDA server, WebSocket protocol, web client). Spirit LM [spiritlm2024] from Meta interleaves speech and text tokens but is research-only. Level 2 (speech-augmented LLMs) models are the largest category. Qwen3-Omni [qwen3omni2025] represents a major advance: it uses an MoE-based Thinker-Talker architecture (30B total, 3B active) with a multi-codebook audio decoder designed for low-latency streaming, supports 119 text languages and 19 speech input languages, and is the first Level 2 model to support function calling. Its predecessor Qwen2.5-Omni [qwen25omni2025] uses a dense 7B Thinker-Talker with TMRoPE for multimodal alignment but lacks function calling and streaming audio. Other Level 2 models include GLM-4-Voice [glm4voice2024] (flow-matching decoder for streaming synthesis), Kimi-Audio [kimiaudio2025] (7B, trained on 13M+ hours), Step-Audio [stepaudio2025] (130B parameters), LLaMA-Omni [llamaomni2024] (226ms latency with Llama-3.1-8B), Mini-Omni [miniomni2024] (“talking while thinking”), and Freeze-Omni [freezeomni2024] (frozen text LLM to prevent catastrophic forgetting). Apart from Qwen3-Omni, none of these models support function calling.
@@
Unlike the frameworks above which provide ready-made components, our contribution is educational: a from-scratch implementation that explains how each piece works internally. While end-to-end models like Qwen3-Omni show great promise, the lack of a self-hostable realtime audio generation solution means the cascaded pipeline remains the practical choice for production deployment. Our tutorial fills this gap by building every component from scratch.
@@ 10 Conclusion
    End-to-end S2S models are approaching realtime, but are not yet self-hostable: Qwen3-Omni achieves $\sim$702ms TTFA and supports function calling—a dramatic improvement over Qwen2.5-Omni ($\sim$13s TTFA, no function calling). However, this latency is only achievable via the cloud-hosted DashScope API. Locally, only the Thinker (text generation) can be served efficiently via vLLM; the Talker (audio synthesis) lacks an optimized self-hosted serving solution.
@@
    The cascaded pipeline is still necessary: For fully self-hosted realtime voice agents with function calling and streaming audio output, the STT $\rightarrow$ LLM $\rightarrow$ TTS cascaded pipeline remains the only viable architecture, achieving $\sim$755ms TTFA.
@@
    Realtime = streaming + pipelining: The perception of instant response comes from overlapping execution of components. The sentence buffer is the critical orchestration primitive.
@@ Appendix A Tutorial Chapter Overview
Appendix B Practical Notes
  1. 1.
    Qwen3-Omni requires vLLM $\geq$0.13.0: The standard vLLM 0.8.x does not recognize Qwen3-Omni as a multimodal model. Version 0.13.0 (or the qwen3_omni branch of vLLM-Omni) is required.
  2. 2.
[…]
  3. 3.
    huggingface-hub version: huggingface-hub>=1.0 breaks transformers==4.57.3. Pin to <1.0.
  4. 4.
    Qwen2.5-Omni transformers version: transformers>=5.0 produces noisy audio with Qwen2.5-Omni. Only ==4.52.3 works correctly.
  5. 5.
    vLLM version for cascaded pipeline: v0.16.0 has a SageMaker dependency that fails outside AWS. v0.8.5 works reliably for Qwen2.5-7B-Instruct.
  6. 6.
  7. 7.
  8. 8.
    Empty SSE chunks: Always guard with if not chunk.choices: continue.
  9. 9.